OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[(x^2+2x-3) / (x^2 + 8x + 16) \div (x-1)/(3x+12)\]

OpenStudy (lynfran):

try factoring if possible first..

OpenStudy (anonymous):

I know you can factor the first side but not sure about the second. Factoring gets me \[(x+3)(x-1) / (x+4)(x+4)\]

OpenStudy (anonymous):

Not sure where to go from there.

OpenStudy (lynfran):

we can (3x+12)= 3(x+4)

OpenStudy (lynfran):

so we have \[\frac{ (x+3)(x-1) }{ (x+4)(x+4) }/\frac{ (x-1) }{ 3(x+4) }\]\[\frac{ (x+3) (x-1)}{ (x+4)(x+4) }*\frac{ 3(x+4) }{ (x-1) }\]

OpenStudy (lynfran):

we can simplify now..

OpenStudy (lynfran):

we have \[\frac{ (x+3) }{ (x+4)}*\frac{ 3 }{ 1 }\]

OpenStudy (anonymous):

That's what I got too. So the answer is 3(x+3) / (x+4)?

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

Thanks!

OpenStudy (lynfran):

welcome